Stockholders’ equity and total assets were $128,000 and $316,000 respectively at the beginning of the period. Assets increased 50% and liabilities decreased 60% during the period.

What is stockholders’ equity at the end of the period?
Answer #1

Total assets=Total liabilities+Total equity

Hence total liabilities at beginning=316,000-128,000=$188,000

Ending assets=316,000*1.5=474,000

total liabilities at end=188,000*(1-0.6)=$75200

Hence ending stockholders equity=474,000-75200

=$398800